OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f o g)(-5) A. -53 B. -9 C. 3 D. 23

OpenStudy (anonymous):

(f o g) f(x)= -2x-7 g(x)==-4x+3 (f o g) = -2(-4x+3)-7 (f o g)= 8x-6-7 substitute -5 to x =8(-5)-6-7 =-40-6-7 =-53